# USC: Uncompromising Spatial Constraints for Safety-Oriented 3D Object Detectors in Autonomous Driving

### v4

@@ IV-C Consolidation for Model Evaluation
As seen, USC is designed as a TP measure to examine a pair of matched prediction $\mathbf{P}$ and ground truth $\mathbf{G}$. After examining the matched pairs for an object class, we can calculate an average USC (AUSC). Then, summarizing the results from all object classes gives the mean AUSC (mAUSC). Furthermore, to also reflect FP and FN predictions, we follow the nuScenes protocol, which incorporates already the recall-precision-based mAP metric, and simply extend mAUSC as follows:
@@
As such, USC-NDS not only reflects the general object detection performance but also characterizes our safety-oriented spatial constraints from the driving perspective. To see their efficacy, we shall use the consolidated AUSC, mAUSC and USC-NDS for model evaluation in our experiments later.
V Safety-Oriented Fine-Tuning
Apart from examining model performance with the safety-oriented spatial constraints, we consider how to improve them via a loss function for continual learning. In particular, along with our analysis, we notice that an ultimate safety-oriented goal in AD is to encapsulate the ground truths with predictions. That is, while the spatial constraints tolerate certain errors and require a prediction to cover the ground truth from the AV’s perspective, we directly aim for proper containment with the 3D bounding boxes during optimization.
It turns out that a proper containment can be modeled by the 3D version of the IoGT function, comparing volumes of the intersection $\mathbf{P}\cap\mathbf{G}$ and ground truth $\mathbf{G}$, similar to Eq. (6). The better a prediction contains the ground truth, the higher its IoGT score. With a full containment, IoGT maximizes at $1$. Therefore, we compose an IoGT loss as:
@@ V Safety-Oriented Fine-Tuning
Nevertheless, the IoGT loss by itself does not provide a complete solution, since it vanishes when the prediction over-approximates the ground truth. We consequently combine it with accuracy-guided loss functions such as the SmoothL1 loss ($L_{\mathsf{SmoothL1}}$[17] and obtain a safety-oriented loss function:
@@
VI Experiments
With the proposed USC-based metrics for model evaluation in Sec. IV-C and the safety-oriented loss function for model fine-tuning in Sec. V, we now conduct experiments to answer the following three research questions (RQs).
@@ VI Experiments
VI-A Model Evaluation
We now start with RQ1 and evaluate various models in terms of the exiting accuracy-based mAP and NDS as well as our USC-based metrics.
@@ VI-A Model Evaluation
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Qualitative fine-tuning results of the camera-based PGD model [19]. We provide samples from two driving scenarios (organized column-wise), each with two time steps. For each time step, the upper left image shows red predictions based on $L_{\mathsf{SmoothL1}}$ [17], the lower left image shows blue predictions based on our $L_{\mathsf{safety}}$, and on the right is a BEV projection of all predictions alongside the ground truths in gray.
TABLE I: Quantitative fine-tuning results of the camera-based PGD model [19].
@@
VI-A1 Implementation
We utilize object detectors hosted on the MMDetection3D platform [22], specifically the top-performing lidar-based models SSN [20] and CenterPoint [21] as well as camera-based ones FCOS3D [18] and PGD [19]. The MMDetection3D platform also has an interface connecting the nuScenes dataset, which defines the evaluation protocol [1] introduced in Sec. III.
Given our safety-oriented considerations for collision avoidance, we make several adaptations on the nuScenes protocol. First, we concentrate on objects within $20$ meters, instead of $50$. Then, we separate the objects by their ranges to give a more detailed analysis. For objects within 00 to $10$ meters, we tighten the TP threshold from $2$ to $1$ meter. Lastly, some object classes do not exist in specific ranges. For example, no construction vehicles are situated within $[0,10]$ meters in all scenes. For such cases, the original nuScenes protocol assumes the worst scores, e.g., 00 AP and full TP errors, leading to overly underestimated performance. Finding that sub-optimal, we do not account for missing classes in our experiments.
VI-A2 Results and Discussions
Second, looking at the plot for $[0,10]$ meters, the mAUSC score of PGD is actually lower than that of FCOS3D, despite it having higher mAP and NDS. A similar situation appears for the lidar-based CenterPoint and SSN. This finding resonates with the related work [10], pointing out that more accurate models may still exhibit safety concerns in nearby ranges.
VI-B Model Fine-Tuning
@@
VI-B1 Implementation
Since the PGD object detector originally outputs a unique image-based representation of the 3D bounding boxes, we append to it a transformation function and obtain the typical representation (as shown in Sec. III). Then, together with the annotated ground truths, the IoGT loss and safety-oriented loss (with $\lambda$ set to $0.8$) can be calculated. We run five fine-tuning sessions for the best PGD checkpoint provided on MMDetection3D [22] using the same learning configurations as the original PGD work, e.g., NN learning rate and data augmentation policy. For comparison, we also run five sessions with the accuracy-guided loss function $L_{\mathsf{SmmoothL1}}$. All sessions are run with a batch size of six samples for six epochs of a subset shared by the nuScenes benchmark [1]. Each session takes roughly two hours on an Nvidia RTX A6000.
VI-B2 Results and Discussions
We present in Tab. I the fine-tuning results, which are the average of the five sessions for the $L_{\mathsf{SmmoothL1}}$-based models and the $L_{\mathsf{safe}}$-based models respectively. We also provide qualitative samples from two driving scenarios, each with two time steps, in Fig 6.
Interestingly, as Tab. I shows, the proposed loss function not only improves the model’s safety-oriented performance in terms of mAUSC but also maintains, or even slightly increases, the accuracy-based NDS. This likely stems from the explicit usage of IoGT in the loss function, which attempts not only to ensure full coverage but to encapsulate the ground truth entirely, thereby keeping the predictions aligned at the same time. The effect is visualized in the qualitative samples, where most of the blue predictions contain the ground truths to a larger extent and are aligned well.
Regarding the class-specific results, a similar trend can be observed too. For example, corresponding to the higher pedestrian AUSC achieved by $L_{\mathsf{safe}}$, the pedestrians are better captured by the blue predictions shown at the lower instance of the right scenario in Fig. 6.
Overall, the proposed loss function does increase the model’s performances, especially in terms of the safety-oriented scores. Although the improvement might be marginal, such differences may play an important role when it comes to safety-critical situations. For further enhancement, nonetheless, one may consider adopting related approaches such as box enlarging in a post-processing fashion [3, 4].
VI-C Closed-Loop Validation
@@
VI-C1 Implementation
@@
To briefly explain, in every simulation, the AV’s mission is to perform valet parking by driving along a designated path from a hand-off zone to a parking slot. During the mission, a pedestrian will cross the driveway, thereby testing the object detection and AEB functions of the AV. The AEB function operates based on a simple rule: If an object is detected within the braking range of the AV, it will take a full brake. Hence, whether a collision occurs depends mainly on the object detector’s performance. Lastly, for more diverse testing, we alter the pedestrian types, speed, crossing timing, and environmental conditions such as illumination. Fig. 7 illustrates the setup. Readers can check our demonstration video [23] or Sec. 4.2 of an overviewing project paper [24] for more details regarding the AV’s simulation and testing.
VI-C2 Results and Discussions
@@
It can be observed that compared to the conventional mAP, NDS already reflects the system-level performance well. This is similarly concluded in the recent validation work [9]. Nonetheless, the proposed mAUSC outperforms NDS, and by combining the two, one can attain the most indicative model-level metric for the system-level collision rates.
To reflect on the results, we find that in most simulations, the pedestrians are detected (as in Fig. 6), and mAUSC successfully characterizes the safety-oriented properties in these cases. Still, there are simulations where the pedestrians are not detected at all, and considering these cases, the well-rounded USC-NDS metric is indeed the most suitable proxy for system-level performance, as shown.
VII Conclusion
In this work, observing the urge for appropriate performance indicators for 3D object detectors, we formulated uncompromising spatial constraints (USC) and corresponding quantitative metrics based on a safety point of view. Concretely, given a prediction $\mathbf{P}$ and its targeted ground truth $\mathbf{G}$, the proposed USC require the prediction $\mathbf{P}$ to cover the ground truth $\mathbf{G}$ when seen from the AV. Such a requirement foregoes the likely unachievable perfect detection demanded by conventional accuracy-based metrics such as mAP. We conducted extensive experiments with the nuScenes datasets and a closed-loop AEB system simulation. The results demonstrated the efficacy of USC-based metrics in finding safety-oriented and more dependable models as well as its potential to gear existing models towards better performances beyond pure accuracy.

### v5

@@ III-C Consolidation for Model Evaluation
As seen, USC is designed as a TP measure to examine a pair of matched prediction $\mathbf{P}$ and ground truth $\mathbf{G}$. After examining the matched pairs for an object class, one can calculate an average USC (AUSC). Then, summarizing the results from all object classes gives the mean AUSC (mAUSC). Furthermore, to also reflect false-positive (FP) and false-negative (FN) predictions, we follow nuScenes’ protocol [2], which incorporates already the recall-precision-based mAP metric, and simply extend mAUSC as follows:
@@
As such, USC-NDS not only reflects the generic object detection performance but also characterizes our safety-oriented spatial constraints from the driving perspective. To see their efficacy, we shall use the consolidated AUSC, mAUSC, and USC-NDS for model evaluation in our experiments later.
IV Safety-Oriented Fine-Tuning
Apart from examining model performance with the safety-oriented spatial constraints, we consider how to improve them via a loss function for continual learning. In particular, along with our analysis, we notice that an ultimate safety-oriented goal in AD is to encapsulate the ground truths with predictions. That is, while the spatial constraints tolerate certain errors and require a prediction to cover the ground truth when seen from the AV, we directly aim for proper containment with the 3D bounding boxes during optimization.
It turns out that a proper containment can be modeled by the 3D version of the IoGT function, comparing volumes of the intersection $\mathbf{P}\cap\mathbf{G}$ and ground truth $\mathbf{G}$, similar to Eq. (3). The better a prediction contains the ground truth, the higher its IoGT score. With full containment, IoGT maximizes at $1$. Therefore, we compose an IoGT loss as:
@@ IV Safety-Oriented Fine-Tuning
Nevertheless, the IoGT loss by itself does not provide a complete solution since it vanishes when the prediction over-approximates the ground truth. We consequently combine it with accuracy-guided loss functions such as the SmoothL1 loss ($L_{\mathsf{SmoothL1}}$[17] and obtain a safety-oriented loss function:
@@
V Experiments
With the proposed USC-based metrics for model evaluation in Sec. III-C and the safety-oriented loss function for model fine-tuning in Sec. IV, we now conduct experiments to answer the following three research questions (RQs).
@@ V Experiments
V-A Model Evaluation
We now start with RQ1 and evaluate various models in terms of the existing accuracy-based mAP and NDS, as well as our USC-based metrics.
@@ V-A Model Evaluation
V-A1 Implementation
As introduced, we perform model evaluation using the nuScenes dataset [2]. Given our safety-oriented considerations for collision avoidance, we make several adaptations to its evaluation protocol. First, we concentrate on objects within $20$ meters instead of $50$. Then, we separate the objects by their ranges to give a more detailed analysis. For objects within 00 to $10$ meters, we tighten the TP threshold for center distances from $2$ to $1$ meters. Lastly, some object classes do not exist in specific ranges. For example, no construction vehicles are situated within $[0,10]$ meters in all scenes. For such cases, the original nuScenes protocol assumes the worst scores, e.g., 00 AP and full TP errors, leading to overly underestimated performance. Finding that sub-optimal, we do not account for missing classes in our experiments.
For the object detection models, we employ the comprehensive MMDetection3D library [22] We adopt the top-performing ones in the nuScenes benchmark [2], including camera-based FCOS3D [18] and PGD [19] as well as lidar-based SSN [20] and CenterPoint [21].
V-A2 Results and Discussions
Second, looking at the plot for $[0,10]$ meters, the mAUSC score of PGD is actually lower than that of FCOS3D, despite it having higher mAP and NDS. A similar situation appears for the lidar-based CenterPoint and SSN. This finding resonates with the related work [11], pointing out that more accurate models may still exhibit safety concerns in short ranges.
V-B Model Fine-Tuning
TABLE I: Quantitative results of fine-tuning the camera-based PGD model [19]. NDS and ATE are accuracy-oriented metrics from nuScenes’ protocol [2], whereas mAUSC and AUSC are our safety-oriented metrics. Using the proposed $L_{\mathsf{safety}}$ not only achieves better safety-oriented scores but also maintains the accuracy.
@@ V-B Model Fine-Tuning
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Quantitative results of fine-tuning camera-based PGD model [19]. For each of the two scenes, the upper left image shows red predictions based on $L_{\mathsf{SmoothL1}}$ [17], the lower left image shows blue predictions based on our $L_{\mathsf{safety}}$, and on the right is a BEV projection of all predictions alongside the ground truths in gray. In general, the blue predictions contain the ground truths to a larger extent while staying aligned with the ground truths.
@@
V-B1 Implementation
Since the PGD object detector originally outputs a unique image-based representation during the learning phase [19], we append to it a transformation function and obtain the 3D bounding box representation. Then, together with the annotated ground truths, the IoGT loss and safety-oriented loss (with $\lambda$ set to $0.8$) can be calculated. We run five fine-tuning sessions for the best PGD checkpoint provided on MMDetection3D [22] using the same learning configurations as the original PGD work, e.g., NN learning rate and data augmentation policy. For comparison, we also run five sessions with the accuracy-guided loss function $L_{\mathsf{SmoothL1}}$. All sessions are run with a batch size of six samples for six epochs of a subset shared by the nuScenes benchmark [2]. Each session takes roughly two hours on an Nvidia RTX A6000.
V-B2 Results and Discussions
We present in Tab. I the fine-tuning results, which are the average of the five sessions for the $L_{\mathsf{SmoothL1}}$-based models and the $L_{\mathsf{safety}}$-based models respectively. We also provide qualitative samples from two scenes in Fig 6.
Interestingly, as Tab. I shows, the proposed loss function not only improves the model’s safety-oriented performance in terms of mAUSC but also maintains, or even slightly increases, the accuracy-based NDS. This likely stems from the explicit usage of 3D IoGT in the loss function defined by Eq. (11), which aims not only at full coverage from the AV’s perspective but at proper containment of the ground truths, thereby keeping the predictions aligned at the same time. The effect is visualized in the qualitative samples, where most of the blue predictions contain the ground truths to a larger extent and are aligned well.
A similar trend can be observed regarding the class-specific results: The safety-oriented models achieve better AUSC scores while maintaining similar average translation error (ATE). For example, corresponding to the higher AUSC, the pedestrians are better contained by the blue predictions in the right scene of Fig. 6.
In summary, the proposed loss function $L_{\mathsf{safety}}$ does improve model performances, especially regarding the safety-oriented scores. Although the improvement might be marginal, such differences may play an important role in safety-critical situations. For further enhancement, nonetheless, one may consider adopting complementary approaches such as box enlarging in a post-processing fashion [4, 5]. Moreover, we leave a full examination and reconciliation of the potentially competing accuracy- and safety-oriented objectives in $L_{\mathsf{safety}}$ as future work.
V-C Closed-Loop Validation
@@
V-C1 Implementation
To briefly explain, in every simulation, the AV’s mission is to perform valet parking by driving along a designated path from a hand-off zone to a parking slot. During the mission, a pedestrian will cross the driveway, thereby testing the object detection and AEB functions of the AV. The AEB function operates based on a simple rule: If an object is detected within the braking range of the AV, it will take a full brake. Hence, whether a collision occurs depends mainly on the object detector’s performance. Lastly, for more diverse testing, we alter the pedestrian type, speed, crossing timing, the AV’s speed, and environmental conditions such as illumination. Fig. 7 illustrates the setup. Readers can check our demonstration video [23] or Sec. 4.2 of an overviewing project paper [24] for more details regarding the AV’s simulation and testing.
@@ V-C1 Implementation
V-C2 Results and Discussions
@@
Compared to the conventional mAP, NDS already reflects the system-level performance well. This is similarly concluded in the recent validation work [10]. Nonetheless, the proposed mAUSC outperforms NDS, and by combining the two, one can attain the most indicative model-level metric for the system-level collision rates.
To reflect on the results, we find that in most simulations, the pedestrians are detected (as in Fig. 6), and mAUSC successfully characterizes the safety-oriented properties in these TP cases. Still, there are cases where the pedestrians are not detected at all or the predictions are too distant from the pedestrians. Considering these cases, the well-rounded USC-NDS metric, which also reflects such FNs and FPs, is indeed the most suitable proxy for system-level performance.
VI Conclusion
In this work, observing the urge for appropriate performance indicators for 3D object detectors, we formulated uncompromising spatial constraints (USC) and corresponding quantitative metrics based on a safety point of view. Concretely, given a prediction $\mathbf{P}$ and its targeted ground truth $\mathbf{G}$, the proposed USC requires the prediction $\mathbf{P}$ to cover the ground truth $\mathbf{G}$ when seen from the AV. Such a requirement foregoes the likely unachievable perfect detection demanded by conventional accuracy-based metrics such as mAP. We conducted extensive experiments with the nuScenes dataset and a closed-loop AEB system simulation. The results demonstrated the efficacy of USC-based metrics in finding safety-oriented and more dependable models as well as its potential to gear existing models towards better performances beyond pure accuracy.